# A Hybrid Strategy for Aggregated Probabilistic Forecasting and Energy Trading in HEFTCom2024

### v1

@@ 3.5 Aggregated Quantile Forecasting
To obtain quantile forecasts for total power generation, it is necessary to derive the CDF of total generation conditioned on covariates $\mathbf{x}$, where $\mathbf{x}$ represents the combined features used for wind and solar power, i.e., $\mathbf{x}=[\mathbf{x}_{\text{w}}^{T},\mathbf{x}_{\text{s}}^{T}]^{T}$. The primary approach involves reconstructing the PDF and CDF for wind and solar power based on their respective quantile forecasts. These reconstructed distributions are then numerically convolved to derive the PDF and CDF of total generation. A key assumption for the numerical convolution of the PDFs for wind and solar power is that they are conditionally independent given the covariates $\mathbf{x}$. This assumption has been validated through regression-based testing [34], with the results provided in the Appendix A. The proposed method is non-parametric and requires the following steps to be implemented.
@@ 3.6 Train-test Framework for Model Tuning
We design a three-fold cross-validation framework to tune the hyperparameters of the LightGBM models. The test set is constructed to ensure consistency with the data quality encountered in online testing. Specifically, the reference time for the test set is fixed at 00:00, with a lead time ranging from 23 to 47 hours. Optuna [35] is used to optimize the hyperparameters of LightGBM. Optuna is an automated hyperparameter optimization framework that searches the hyperparameter space in parallel. Specifically, the type of hyperparameter and the range of values are shown in Table 1. The final hyperparameters are selected based on their performance on the three-fold cross-validation.
@@ 4.1 Analysis of Electricity Prices
Figure 5: The mean and variability of the price spread at different hour-of-day in various periods. Each subplot represents the statistical data of a half-year period. The x-axis represents different hours of the day, and the y-axis represents the mean and standard deviation of the price spread.
@@ 4.2 Stochastic Trading Model
$\max_{\hat{e}_{1},\dots,\hat{e}_{48}}\mathbb{E}[S_{trading}]=\mathbb{E}[\sum_{t=1}^{48}\pi_{\text{D},t}\hat{e}_{t}-0.07(\hat{y}_{t}-\hat{e}_{t})^{2}]+\sum_{t=1}^{48}\pi_{\text{SS},t}y_{t},$ (22)
$s.t\quad 0\leq\hat{e}_{t}\leq 1800\quad\forall t=1,2,...,48,$
@@
$\hat{\mathbb{E}}(\pi_{\text{D},t}|t)=\frac{1}{\left\lfloor w/48\right\rfloor}\sum_{i=1}^{w}\pi_{\text{D},i}\mathbb{I}(i\bmod 48=t),$ (24)
@@
$\max_{\hat{e}_{t}\in[0,1800]}r(\hat{e}_{t})=\hat{\mathbb{E}}(\pi_{\text{D},t}|t)\hat{e}_{t}-0.07(\hat{y}_{t}-\hat{e}_{t})^{2}.$ (25)
Given the power forecast $\hat{y}_{t}$ and the expected estimate of the price spread $\hat{\mathbb{E}}(\pi_{\text{D},t}|t)$, the optimal bidding decision $\hat{e}^{*}_{t}$ can be obtained through the Karush-Kuhn-Tucker (KKT) conditions of Eq. (25). The optimal bidding strategy is given by
@@ 4.3 Power Forecasting Model for Trading
$\mathcal{L}(\hat{y}_{t},\hat{\pi}_{\text{D},t},y_{t},\pi_{\text{D},t})=0.07(\hat{y}_{t}-y_{t})^{2}+3.57(\hat{\mathbb{E}}(\pi_{\text{D},t}|t)-\pi_{\text{D},t})^{2}+(\hat{y}_{t}-y_{t})(\hat{\mathbb{E}}(\pi_{\text{D},t}|t)-\pi_{\text{D},t}).$ (28)
@@
[…]
@@ 5.1 Evaluation Environments and Metrics
[…]
where $\hat{F}(y)$ is the CDF of the probabilistic forecast, and $y_{i}$ is the actual power generation. Winkler Score is a metric that evaluates the accuracy and width of the forecast. It penalizes both the accuracy and width of the forecast, defined as
@@ 5.2 Evaluation of Stacking Sister Models
DWD 28.96 53.20 358.12 18.88 34.64 231.00 13.34 24.28 164.13 13.71 24.92 156.46
GFS 30.43 55.89 380.96 19.22 35.31 244.03 17.52 31.99 217.74 15.04 27.36 170.77
Stacking 27.13 49.73 334.58 17.70 32.47 220.21 13.33 24.29 163.80 13.66 24.86 155.34
In the wind power forecasting task, the stacked models significantly reduce the MPL, MCRPS, and MWS compared to models trained on a single NWP in both cases. For the MPL metric, which is the focus of the competition, the stacked models achieve improvements of 10.84% and 6.32% compared to the GFS and DWD models in Case I. In Case II, the stacked models achieve improvements of 7.91% and 6.25% compared to the GFS and DWD models, respectively. However, for the solar power forecasting task, the effect of the stacked models is minimal, only reducing the MPL by 0.075% and 0.36% in Case I and Case II, respectively, compared to using only the NWP provided by DWD. Even in Case I, the MCRPS of the stacked models is slightly higher than that of the models trained on a single NWP. The detailed reason for this result is explained in Section 6.
@@ 5.3 Evaluation of Online Post-processing
Offline Model 15.25 27.25 188.97
Offline Post-processing 14.06 25.51 157.66
Online Post-processing 13.71 24.92 156.46
The rolling test results show that compared to the model trained solely on the offline dataset, the MPL, MCRPS, and MWS are reduced by 10.1%, 8.55%, and 17.2%, respectively, and compared to the offline post-processing model, the MPL, MCRPS, and MWS are also reduced by 2.49%, 2.31%, and 0.76%, respectively. This also indicates that the online post-processing approach is more capable of dynamically adapting to the changed distribution of solar power generation. Figure 6 illustrates the quantile forecast results alongside the actual solar power values. Despite the reduced sharpness on certain days, notably, the original forecasts failed to capture the peak solar power values on certain days, (e.g., 03-06, 03-08, 03-09). The proposed online rolling post-processing model overcomes this limitation and achieves better results in MCRPS and MWS.
@@ 5.4 Evaluation of Aggregated Quantile Forecasting
Quantile by Quantile 34.41 47.99 417.33 24.48 37.91 290.56
Total Forecasting 34.84 43.38 428.46 none none none
Aggregation 34.36 46.88 418.37 24.29 37.14 285.99
@@
For the MPL metric, the proposed quantile aggregation method reduces the MPL compared to the summation method quantile by quantile in both cases, with reductions of 0.15% and 0.78%, respectively. Compared to directly forecasting the total generation, the proposed method also achieves superior performance with a 1.38% reduction. For the MCRPS metric, the proposed method shows a more significant reduction compared to the MPL metric, with reductions of 2.31% and 2.03%, respectively. This reflects the predicted CDF from the proposed method is more rational than the quantile-by-quantile method. Although the direct total generation forecasting method achieves best fit to the actual CDF (i.e., the lowest MCRPS), it does not achieve better performance in the MPL metric, which is the focus of the competition. The increased input dimensionality required for modeling the aggregated output may lead to a greater risk of overfitting and makes it more challenging to design specialized hyperparameters for individual quantile forecasting models. The recorded performance of the GEB team is also shown in Table 4, which reflects the actual scores obtained during the HEFTCom2024 competition. It is observed that the proposed method slightly outperforms the GEB team during the competition. This is because our method was gradually developed during the competition, while the case study used the final solution.
@@ 5.5 Evaluation in Trading Tasks
Bid as MSE-oriented 306.26 0.10 87.63 0.01
Naive Forecasting 305.54 –0.13 87.51 –0.13
Seasonal Persistence Model 306.03 0.03 87.64 0.02
AR 306.18 0.08 87.62 0.00
ETS 306.38 0.14 87.76 0.16
ARMA 306.19 0.08 87.62 0.00
SARMAX 306.20 0.08 87.63 0.01
ST (q50) 307.22 0.42 88.00 0.43
ST (MSE-oriented) 307.26 0.43 88.00 0.43
Table 5 summarizes the revenue performance of various trading strategies across two cases. Although a variety of spread forecasting methods are compared to provide informative signals for trading, the proposed trading strategy consistently achieved the best overall performance in both cases, with the improvement of roughly 0.44% over the “bid as forecast” baselines. Besides, the table reveals several noteworthy observations.
For the power generation forecast required by the trading strategy, Simply changing the loss function of the power forecasting model from the pinball loss at 50% quantile to MSE can achieve a modest revenue improvement (e.g., from £305.94 m to £306.26 m in Case I), as MSE is more similar to the downstream trading loss structure.
@@ 6.1 Tricks and Lessons
In general, we believe that identifying the key issues in the HEFTCom2024 forecasting and trading tasks and designing tailored methods is crucial to achieving high performance. We tend to choose a combination of multiple lightweight and well-established methods to address these specific problems. This allows for developing each module independently. For the first issue, results of case studies show a reduction in pinball loss when combining multi-source NWP data in wind power forecasting. The stacking technique provides a flexible and robust ensemble method. However, the effectiveness of stacking hinges on combining strong yet diverse base learners. In our view, the discrepancies in forecast performance among sister models are largely attributable to variations in the spatial coverage of the NWP data. As illustrated in Section 2.3, for the Hornsea 1 area, the NWP coordinates of DWD and GFS do not completely overlap. Although predictions solely from GFS underperform compared to DWD, their combination via stacking actually yields superior results. Conversely, in solar power forecasting, the NWP data share identical coverage areas, meaning that the stacking essentially combines only the underlying NWP model differences, which likely leads to a minimal impact. For the second issue, the change in installed capacity results in a new distribution of solar data. Directly correcting the power forecasts based on the proportion of capacity changes is unreliable, as the growth in capacity is not spatially uniform, and the capacity data is not perfect. It is more reliable to address the distribution shift through data-driven methods, such as learning an online post-processing model. This method can also be applied to other online forecasting scenarios with a deterministic and simple distribution shift. Third, for tasks that require predicting aggregate totals, a decomposition-based forecast is generally appropriate, as it allows for targeted handling of individual components. The proposed method offers a theoretically seamless trick for summing individual probabilistic forecasts and is adaptable to various models capable of learning the pinball loss, although it is computationally expensive as a sufficient number of models need to be trained to fit the CDF. However, given that sufficient offline preparation time is available, the computational cost associated with enhancing accuracy is of relatively minor concern. In our final solution, employing 101 quantile levels is found to achieve an acceptable trade-off between predictive resolution of CDF and training efficiency, with the complete training process requiring approximately 2–3 hours. While increasing the number of quantiles enables a finer reconstruction of the CDF, however, due to the multiplicative nature of the model configurations involved, the resulting increase in training time may hinder overall development efficiency. Besides, due to different forecasting tasks and resource constraints, it is advisable to conduct a case-by-case analysis to determine the most suitable number of quantile models. Fourth, although the price spread is a stationary series signal-to-noise ratio, its seasonal characteristics still provide valuable information for enhancing trading performance. The proposed ST method can achieve higher revenue compared to “bid as power forecast” and AR, MA or persistence based models, as it is less sensitive to noise.
@@ 6.2 Further Improvements to Trading Strategies
To validate the effectiveness of the proposed improvement, we conducted detailed experiments on the Case I and Case II datasets. A multilayer perceptron (MLP) is employed as the E2E trading model, as it allows for flexible loss function design and optimizing the trading loss through differentiable backpropagation. The input features consist of the total power forecast, one-hot encoded hour-of-day. Additionally, an extra benchmark is introduced, which disregards the surface shape of trading loss with respect to the two error components, denoted as “accuracy-oriented” in Table 6. This benchmark shares identical input features with the proposed E2E trading model but adopts MSE as the loss function to predict the price spread. The optimal bidding strategy is derived based on the KKT conditions. The test results on both cases are presented in Table 6. It can be observed that the proposed E2E trading model achieves the highest revenue in both cases. Compared to the “bid as q50” and ST (MSE-oriented) methods, the revenue increases by 0.59% and 0.16% in Case I, and by 0.57% and 0.14% in Case II, respectively. While the absolute percentage gains may appear limited, they are nevertheless notable given the low signal-to-noise ratio inherent in the price spread input.
@@
bid as q50 305.94 82.92 87.62 16.92
ST(final solution) 307.26 81.61 88.00 16.53
Accuracy-oriented 306.37 82.50 87.76 16.77
E2E 307.75 81.11 88.12 16.41
To further demonstrate how the E2E approach enhances trading performance by shaping the error distribution, we also visualize the power forecast errors and price spread forecast errors of the compared methods in the scatter plot shown in Figure 9. Compared to the bid-as-q50 method (i.e., the estimated price spread is always zero), the ST (final solution of GEB) and accuracy-oriented methods primarily improve trading revenue mainly through more precise estimation of the price spread, which results in an error distribution that is more tightly clustered around zero. Notably, the error distribution of the E2E approach is better aligned with the shape of the trading loss surface depicted in Figure 8. To highlight the fundamental differences between the proposed method and others, we emphasize a subset of points in the subplot of E2E approach. These points exhibit higher revenue compared to the ST method but also have higher forecast errors. We refer to these points as ”Sample of Decision-focused Revenue (SoDR)”. The SoDR points tend to be distributed with a certain tilt and are primarily located in regions associated with low trading loss, while appearing less frequently in high-loss regions. Overall, since the trading model is fixed as a deterministic algorithm, the trading revenue is mainly determined by the quality of the forecasts. Even with imperfect predictions, the trading revenue can be improved by training forecasts that are closer to the downstream decision-making loss.
Figure 9: Scatter plot of forecast errors of power generation and price spread for different methods.

### v2

@@ 3.5 Aggregated Quantile Forecasting
To obtain quantile forecasts for total power generation, it is necessary to derive the CDF of total generation conditioned on covariates $\mathbf{x}$, where $\mathbf{x}$ represents the combined features used for wind and solar power, i.e., $\mathbf{x}=[\mathbf{x}_{\text{w}}^{T},\mathbf{x}_{\text{s}}^{T}]^{T}$. The primary approach involves reconstructing the PDF and CDF for wind and solar power based on their respective quantile forecasts. These reconstructed distributions are then numerically convolved to derive the PDF and CDF of total generation. A key assumption for the numerical convolution of the PDFs for wind and solar power is that they are conditionally independent given the covariates $\mathbf{x}$. This assumption has been validated through regression-based testing (Zhang et at.,, 2017), with the results provided in the Appendix A. The proposed method is non-parametric and requires the following steps to be implemented.
@@ 3.6 Train-test Framework for Model Tuning
We design a three-fold cross-validation framework to tune the hyperparameters of the LightGBM models. The test set is constructed to ensure consistency with the data quality encountered in online testing. Specifically, the reference time for the test set is fixed at 00:00, with a lead time ranging from 23 to 47 hours. Optuna (Akiba et al.,, 2019) is used to optimize the hyperparameters of LightGBM. Optuna is an automated hyperparameter optimization framework that searches the hyperparameter space in parallel. Specifically, the type of hyperparameter and the range of values are shown in Table 1. The final hyperparameters are selected based on their performance on the three-fold cross-validation.
@@ 4.1 Analysis of Electricity Prices
Figure 5: The mean and variability of the price spread at different hour-of-day in various periods. Each subplot represents the statistical data of a half-year period. The x-axis represents different hours of the day. The markers represent the mean of the price spread at each hour, while the bars represent the standard deviation of the price spread at each hour.
@@ 4.2 Stochastic Trading Model
[…]
@@
[…]
@@
[…]
Given the power forecast $\hat{y}_{t}$ and the conditional expectation of the price spread $\hat{\mathbb{E}}(\pi_{\text{D},t}|t)$, the optimal bidding decision $\hat{e}^{*}_{t}$ can be obtained through the Karush-Kuhn-Tucker (KKT) conditions of Eq. (25). The optimal bidding strategy is given by
@@ 4.3 Power Forecasting Model for Trading
[…]
@@
$\theta^{*}=\underset{\theta}{\text{argmin}}\quad\frac{1}{|\mathcal{W}_{\text{DWD}}|}\sum_{t=1}^{|\mathcal{W}_{\text{DWD}}|}(F^{\prime}(\mathbf{x}_{t};\theta)-y_{t})^{2}.$ (29)
@@ 5.1 Evaluation Environments and Metrics
$\mathrm{MCRPS}=\frac{1}{N}\sum_{i=1}^{N}\int_{-\infty}^{\infty}(\hat{F}_{i}(y)-\mathbb{I}(y\geq y_{i}))^{2}dy,$ (30)
where $\hat{F}_{i}(y)$ is the CDF of the probabilistic forecast, and $y_{i}$ is the actual power generation. Winkler Score is a metric that evaluates the accuracy and width of the forecast. It penalizes both the accuracy and width of the forecast, defined as
@@ 5.2 Evaluation of Stacking Sister Models
DWD 28.96 53.18 357.72 18.90 34.66 230.98 13.35 24.25 164.28 13.72 24.91 157.30
GFS 30.44 55.87 381.51 19.19 35.27 243.49 17.54 32.01 218.20 15.07 27.35 171.54
Stacking 27.13 49.74 334.65 17.69 32.46 220.20 13.33 24.25 163.84 13.68 24.86 156.06
In the wind power forecasting task, the stacked models significantly reduce the MPL, MCRPS, and MWS compared to models trained on a single NWP in both cases. For the MPL metric, which is the focus of the competition, the stacked models achieve improvements of 10.87% and 6.32% compared to the GFS and DWD models in Case I. In Case II, the stacked models achieve improvements of 7.82% and 6.40% compared to the GFS and DWD models, respectively. However, for the solar power forecasting task, the effect of the stacked models is minimal, only reducing the MPL by 0.15% and 0.29% in Case I and Case II, respectively, compared to using only the NWP provided by DWD. Even in Case I, the MCRPS of the stacked models is slightly higher than that of the models trained on a single NWP. The detailed reason for this result is explained in Section 6.
@@ 5.3 Evaluation of Online Post-processing
Offline Model 15.22 27.18 188.16
Offline Post-processing 14.22 25.85 161.00
Online Post-processing 13.72 24.92 157.30
The rolling test results show that compared to the model trained solely on the offline dataset, the MPL, MCRPS, and MWS are reduced by 9.86%, 8.31%, and 16.4%, respectively, and compared to the offline post-processing model, the MPL, MCRPS, and MWS are also reduced by 3.52%, 3.60%, and 2.3%, respectively. This also indicates that the online post-processing approach is more capable of dynamically adapting to the changed distribution of solar power generation. Figure 6 illustrates the quantile forecast results alongside the actual solar power values. Despite the reduced sharpness on certain days, notably, the original forecasts failed to capture the peak solar power values on certain days, (e.g., 03-06, 03-08, 03-09). The proposed online rolling post-processing model overcomes this limitation and achieves better results in MCRPS and MWS.
@@ 5.4 Evaluation of Aggregated Quantile Forecasting
Quantile by Quantile 34.40 47.99 417.37 24.50 37.77 292.46
Total Forecasting 34.86 43.41 429.19 none none none
Aggregation 34.35 46.88 418.13 24.32 36.93 288.33
@@
For the MPL metric, the proposed quantile aggregation method reduces the MPL compared to the summation method quantile by quantile in both cases, with reductions of 0.15% and 0.73%, respectively. Compared to directly forecasting the total generation, the proposed method also achieves superior performance with a 1.46% reduction. For the MCRPS metric, the proposed method shows a more significant reduction compared to the MPL metric, with reductions of 2.31% and 2.22%, respectively. This reflects the predicted CDF from the proposed method is more rational than the quantile-by-quantile method. Although the direct total generation forecasting method achieves best fit to the actual CDF (i.e., the lowest MCRPS), it does not achieve better performance in the MPL metric, which is the focus of the competition. The increased input dimensionality required for modeling the aggregated output may lead to a greater risk of overfitting and makes it more challenging to design specialized hyperparameters for individual quantile forecasting models. The recorded performance of the GEB team is also shown in Table 4, which reflects the actual scores obtained during the HEFTCom2024 competition. It is observed that the proposed method slightly outperforms the GEB team during the competition. This is because our method was gradually developed during the competition, while the case study used the final solution.
@@ 5.5 Evaluation in Trading Tasks
Bid as MSE-oriented 306.32 0.12 87.68 0.07
Naive Forecasting 305.61 –0.11 87.56 –0.07
Seasonal Persistence Model 306.10 0.05 87.68 0.07
AR 306.24 0.10 87.67 0.06
ETS 306.43 0.16 87.81 0.22
ARMA 306.25 0.10 87.67 0.06
SARMAX 306.26 0.10 87.68 0.07
ST (q50) 307.29 0.44 88.02 0.46
ST (MSE-oriented) 307.39 0.47 88.05 0.49
Table 5 summarizes the revenue performance of various trading strategies across two cases. Although a variety of spread forecasting methods are compared to provide informative signals for trading, the proposed trading strategy consistently achieved the best overall performance in both cases, with the improvement of roughly 0.47%-0.49% over the “bid as forecast” baselines. Besides, the table reveals several noteworthy observations.
For the power generation forecast required by the trading strategy, Simply changing the loss function of the power forecasting model from the pinball loss at 50% quantile to MSE can achieve a modest revenue improvement (e.g., from £305.94 m to £306.32 m in Case I), as MSE is more similar to the downstream trading loss structure.
@@ 6.1 Tricks and Lessons
In general, we believe that identifying the key issues in the HEFTCom2024 forecasting and trading tasks and designing tailored methods is crucial to achieving high performance. We tend to choose a combination of multiple lightweight and well-established methods to address these specific problems. This allows for developing each module independently. For the first issue, results of case studies show a reduction in pinball loss when combining multi-source NWP data in wind power forecasting. The stacking technique provides a flexible and robust ensemble method. However, the effectiveness of stacking hinges on combining strong yet diverse base learners. In our view, the discrepancies in forecast performance among sister models are largely attributable to variations in the spatial coverage of the NWP data. As illustrated in Section 2.3, for the Hornsea 1 area, the NWP coordinates of DWD and GFS do not completely overlap. Although predictions solely from GFS underperform compared to DWD, their combination via stacking actually yields superior results. Conversely, in solar power forecasting, the NWP data share identical coverage areas, meaning that the stacking essentially combines only the underlying NWP model differences, which likely leads to a minimal impact. For the second issue, the change in installed capacity results in a new distribution of solar data. Directly correcting the power forecasts based on the proportion of capacity changes is unreliable, as the growth in capacity is not spatially uniform, and the capacity data is not perfect. It is more reliable to address the distribution shift through data-driven methods, such as learning an online post-processing model. This method can also be applied to other online forecasting scenarios with a deterministic and simple distribution shift. Third, for tasks that require predicting aggregate totals, a decomposition-based forecast is generally appropriate, as it allows for targeted handling of individual components. The proposed method offers a theoretically seamless trick for summing individual probabilistic forecasts and is adaptable to various models capable of learning the pinball loss, although it is computationally expensive as a sufficient number of models need to be trained to fit the CDF. However, given that sufficient offline preparation time is available, the computational cost associated with enhancing accuracy is of relatively minor concern. In our final solution, employing 101 quantile levels is found to achieve an acceptable trade-off between predictive resolution of CDF and training efficiency, with the complete training process requiring approximately 4-5 hours. While increasing the number of quantiles enables a finer reconstruction of the CDF, however, due to the multiplicative nature of the model configurations involved, the resulting increase in training time may hinder overall development efficiency. Besides, due to different forecasting tasks and resource constraints, it is advisable to conduct a case-by-case analysis to determine the most suitable number of quantile models. Fourth, although the price spread is a stationary series with low signal-to-noise ratio, its seasonal characteristics still provide valuable information for enhancing trading performance. The proposed ST method can achieve higher revenue compared to “bid as power forecast” and AR, MA or persistence based models, as it is less sensitive to noise.
@@ 6.2 Further Improvements to Trading Strategies
To validate the effectiveness of the proposed improvement, we conducted detailed experiments on the Case I and Case II datasets. A multilayer perceptron (MLP) is employed as the E2E trading model, as it allows for flexible loss function design and optimizing the trading loss through differentiable backpropagation. The input features consist of the total power forecast, one-hot encoded hour-of-day. Additionally, an extra benchmark is introduced, which disregards the surface shape of trading loss with respect to the two error components, denoted as “accuracy-oriented” in Table 6. This benchmark shares identical input features with the proposed E2E trading model but adopts MSE as the loss function to predict the price spread. The optimal bidding strategy is derived based on the KKT conditions. The test results on both cases are presented in Table 6. It can be observed that the proposed E2E trading model achieves the highest revenue in both cases. Compared to the “bid as q50” and ST (MSE-oriented) methods, the revenue increases by 0.62% and 0.15% in Case I, and by 0.64% and 0.15% in Case II, respectively. Although the absolute percentage gains may appear limited, they carry practical significance in view of the low signal-to-noise ratio of the price spread. Notably, the ST-based strategy improved revenue by just 0.49% over the baseline, but this modest advantage translated into a 3rd-place finish among all teams in HEFTCom2024.
@@
bid as q50 305.94 82.92 87.62 16.91
ST(final solution) 307.38 81.47 88.05 16.48
Accuracy-oriented 306.43 82.43 87.82 16.71
E2E 307.84 81.02 88.18 16.35
To further demonstrate how the E2E approach enhances trading performance by shaping the error distribution, we also visualize the power forecast errors and price spread forecast errors of the compared methods in the scatter plot shown in Figure 9. Compared to the bid-as-q50 method (i.e., the estimated price spread is always zero), the ST (final solution of GEB) and accuracy-oriented methods primarily improve trading revenue mainly through more precise estimation of the price spread, which results in an error distribution that is more tightly clustered around zero. Notably, the error distribution of the E2E approach is better aligned with the shape of the trading loss surface depicted in Figure 8. To highlight the differences between the proposed method and others, we emphasize a subset of points in the subplot of E2E approach. These points exhibit higher revenue compared to the ST method but also have higher forecast errors. They tend to be distributed with a certain tilt and are primarily located in regions associated with low trading loss, while appearing less frequently in high-loss regions. Overall, since the trading model is fixed as a deterministic algorithm, the trading revenue is mainly determined by the quality of the forecasts. Even with imperfect predictions, the trading revenue can be improved by training forecasts that are closer to the downstream decision-making loss.
Figure 9: Scatter plots of forecast errors in power generation and price spread under different forecasting methods. In the subplot which denotes the E2E approach, the highlighted pink points indicate cases where the method achieves higher trading revenue compared to the ST method despite having larger forecast errors, illustrating the benefit of aligning forecasts with decision-making objectives.